A partially dominant mutant allele (R) reduces the number of offspring per mosquito from 600 to 200 for heterozygotes (Rr) and from 600 to 0 for homozygotes (RR). What is the average relative fitness of a population in which the gene frequency of R is 0.30?
Answer #1

Hello,
If R = 0.3, then, by Hardy Weinberg, r = 0.7

How to calculate relative fitness?
Relative fitness is calculated by comparing survival rate/ or reprouctive rate. If a population has only reproductive rate varying, then we divide the reproductive rate of given phenotype to maximal reproductive rate.

Thus relative fitness of different phenotypes/and genotypes are :
rr - 600/600= 1
Rr - 200/600 = 0.33
RR - 0/600 = 0


Now, if R = 0.3 - frequency of RR = (0.3)² (By hardy Weinberg)
=0.09 - These individuals have relative fitness = 0

If r = 0.7 - frequency of rr = (0.7)²
= 0.49 - These individuals have repative fitness = 1

If r= 0.7, R = 0.3
Frequency of Rr = 2 (0.7) (0.3) By hardy weinberg
= 0.42 - These individuals have relative fitness = 0.33

Mean relative fitness of population = [0 + 0.49 + {(0.42)(0.33)}]/3
= 0.629/3 = 0.21


The average relative fitness of entire population is 0.21

Thank you
Hope I helped
Please ask any doubt in comments.